OpenStudy (anonymous):

Solve for x: one over eight (8x + 15) = 24 @SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{1}{8}(8x+5)=24 }\) like this?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

multiply both sides times 8.

OpenStudy (anonymous):

Solve for x: one over eight (8x + 15) = 24 9 49 one hundred seventy-seven over eight two hundred seven over eight

OpenStudy (solomonzelman):

no you didn't multiply the 24 times 8 too.

OpenStudy (solomonzelman):

you got the fact that 1/8 is cancelled, but you didn't multipyl 24 times 8.

OpenStudy (solomonzelman):

\(\large\color{black}{ 8x+5=192 }\)

OpenStudy (solomonzelman):

subtract 5 from both sides, and divide both sides by 8.

OpenStudy (solomonzelman):

you got it, sure?

OpenStudy (solomonzelman):

or didn't get it?

OpenStudy (solomonzelman):

ohh it is 15, not 5.

OpenStudy (solomonzelman):

So it is 8x+15=192.

OpenStudy (solomonzelman):

8x+15=192. 8x=177 x=177/8 or 22.125

OpenStudy (anonymous):

so what is it

OpenStudy (solomonzelman):

C

OpenStudy (anonymous):

ok thanks

OpenStudy (solomonzelman):

yw

OpenStudy (anonymous):

207\[207 \over 8\]